find the slope of a line parallel to 2x+5y=3

Answer by jim_thompson5910(35256)

2x+5y=3

5y=3-2x

5y=-2x+3

y=-2/5x+3/5

So the slope is m = -2/5
--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------